NPP Projection Report 9/4 Admin


To do this evening/ was completed:

Presentation Issues:

Preshow Clips/Playlist:

Trailer Updates:

Special Events/ AV set up: Spoons is all set up for tonight. Wireless mics are set on stands and tested. Yamaha mixer and spotlight controls are on the rollout table with the AJA extender. Spotlights and HDMI are working. I left everything on except the mics, Yamaha mixer, and spotlight control board.

I don't know who broke down the PA system in T1 last so I'm noting this here: please remember to hang the XLR cable for the left speaker on the hook under the curtain, please don't leave it on the floor. The hook is way over to the left under the curtain, attached to the frame of the screen like the hook for the right speaker cable.

Had a walkthrough with Devin from TEAL and Chenelle for the event on Saturday. We went through the run of show for each theater. I left a PA speaker set up in two of the three theaters they've rented so they're ready for Saturday. I'll be running this event.

Ordering/Receiving (posters & trailers):

Ordering/Receiving (content): Jackass print arrived, it's in booth 1. Please inspect this tonight so we can look at it tomorrow.

The download of the Lord of the Rings DCP seems to have failed, Screenwriter was saying it's corrupt and wouldn't transfer to T1, so I restarted the download. It finished successfully and is ingesting.

Desmond sent over a file of Sisterhood of the Traveling Pants which shows for brunch in two weeks. I downloaded it and started converting it to DCP in T1 so we can let him know how it looks on screen sooner rather than later.

Outgoing/Shipping (content): Blonde Venus was picked up.

Fantastic Planet return labels printed and attached to the cans. Pickup is scheduled for tomorrow and the print is on the cart. Emailed tracking info to Brian from Janus.

Print/video Inspection:

Asset tracking doc updated/checked?:

Maintenance/cleaning: Cable connection fixed in Trees. I reseated the ethernet cable behind the A/V bench (under the 2001 70mm strips).

Cue sheets updated/checked (every shift):

Marquee updated?:

Posters updated?:

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

Projection Admin 10/2

Projection Admin 9/18

Projection Admin 9/2